OpenStudy (anonymous):

If the equation y = x + 4 is changed to y = x - 2, how will the graph of the equation change? A) The original graph will shift 2 units down. B) The original graph will shift 6 units down. C) The original graph will shift 6 units to the left. D) The original graph will shift 2 units to the left.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

14) x -5 -3 4 1 -1 -2 0 2 3 -4 y -10 -8 9 1 -2 -6 -1 3 6 -8 What is the equation of the line of best fit for this data? A) y = 2.097x - 0.552 B) y = 2.097x + 0.552 Eliminate C) y = -0.552x + 2.097 D) y = 0.522x - 2.097